OpenStudy (anonymous):

A triangle has sides of 3,6,8, In a similar triangle the shortest side is 9 and the longest is x. Write an proportion of the models?

OpenStudy (anonymous):

you could use \(\frac{8}{3}=\frac{x}{9}\)

OpenStudy (anonymous):

what is the sides were 3,8,10? and shortest side was 6

OpenStudy (anonymous):

would it be 10/3=x/6?